OpenStudy (anonymous):

can someone please help me? i need to simplify this radical... square root of 15a^8 b

OpenStudy (amistre64):

take out all the perfect squares

OpenStudy (amistre64):

a^4 sqrt(15b)

OpenStudy (anonymous):

dont you first have to find out what perfect square number times a regular number wud give you 15?

OpenStudy (anonymous):

like there is no perfect square that can be multiplied to equal 15.

OpenStudy (anonymous):

\[\sqrt{15a ^8b}\]

OpenStudy (amistre64):

yes; 15 has no perfect square factors :)

OpenStudy (anonymous):

Okay. so what can i do?

OpenStudy (amistre64):

8/2 = 4 so pull out a^4

OpenStudy (amistre64):

sqrt(a^8) means a^8/2 :)

OpenStudy (anonymous):

so ull end up with \[\sqrt{15a ^{4}}b\]

OpenStudy (amistre64):

pull out the a's sqrt(a^8) = a^8/2 = a^4 a^4^2 = aa^8

OpenStudy (amistre64):

\[a^4\sqrt{15b}\]

OpenStudy (anonymous):

i got you.

OpenStudy (anonymous):

so that wud be you answer.

OpenStudy (anonymous):

your***?

OpenStudy (amistre64):

it would; yes

OpenStudy (anonymous):

thank you. so much.

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

wait. one question. just wanted to be sure... \[\sqrt[2]{1000}\] is 100 right?